# P.O. Box 15500, 00076 Aalto, Finland

### v1

@@ Appendix A Appendix
One possible interpretation of a shortest path in a graph which has probabilities as edge weights is to interpret the path with the highest probability as if it were the shortest path. If we denote $p_{1},p_{2},\dots,p_{m}$ as the edge probabilities in a graph $G=(V,E)$, where $V$ is the set of nodes, $E$ is the set of edges between nodes, and $m=|E|$ is the number of edges, then our task is to find a path between source node $s$ and target node $t$ that maximizes the product $p_{(s,a)}\cdots p_{(b,t)}$, where $(s,a),\dots,(b,t)\in E$ are the edges and $s,a,\ldots,b,t\in V$ are the nodes on the path that maximizes the probability of influence between nodes $s$ and $t$.
Maximizing the product $p_{(s,a)}\cdots p_{(b,t)}$ is equivalent to maximizing the logarithmic sum logp_(s, a) + …+ logp_(b, t) which is equivalent to minimizing -logp_(s, a) - …- logp_(b, t) which is again equivalent to minimizing log1p(s, a) + …+ log1p(b, t). In the previous formulation, we assumed that the shortest path contains at least one intermediary node. If that is not the case, then naturally the shortest path between $s$ and $t$ is the edge $(s,t)$.
When $0<p_{i}\leq 1$ for all $1\leq i\leq m$, then $\log(1/p_{i})\geq 0$ for all $1\leq i\leq m$. Now, $\log(1/p_{i})$, for all $1\leq i\leq m$, can be used as a nonnegative distance between nodes, because as probability $p$ increases the distance $\log(1/p)$ decreases, i.e. $\log(1/p)$ is a strictly decreasing function of $p$. Furthermore, we can now use Dijkstra’s algorithm for finding shortest paths in the network, since edge weights are non-negative [thomas2009]. In this case, Dijkstra’s algorithm picks the edges with the lowest distance $\log(1/p_{i})$. Therefore, with this transformation the path $i$ with the lowest value $\log(1/p_{i})$ is chosen i.e. the highest probability path is chosen for the shortest path. In the special case, where $p_{i}=1$ for all $1\leq i\leq m$, then the problem reduces to finding the path with the fewest edges.

### v2

@@ Appendix A Appendix
One possible interpretation of a shortest path in a graph that has probabilities as edge weights is to interpret the path with the highest probability as if it were the shortest path. If we denote $p_{1},p_{2},\dots,p_{m}$ as the edge probabilities in a graph $G=(V,E)$, where $V$ is the set of nodes, $E$ is the set of edges between nodes, and $m=|E|$ is the number of edges, then our task is to find a path between source node $s$ and target node $t$ that maximizes the product $p_{(s,a)}\cdots p_{(b,t)}$, where $(s,a),\dots,(b,t)\in E$ are the edges and $s,a,\ldots,b,t\in V$ are the nodes on the path that maximizes the probability of influence between nodes $s$ and $t$.
Maximizing the product $p_{(s,a)}\cdots p_{(b,t)}$ is equivalent to maximizing the sum $\log p_{(s,a)}+\ldots+\log p_{(b,t)}$, which is equivalent to minimizing $\log p_{(s,a)}-\ldots-\log p_{(b,t)}$, which is again equivalent to minimizing $\log(1/p_{(s,a)})+\ldots+\log(1/p_{(b,t)}).$ In the previous formulation, we assumed that the shortest path contains at least one intermediary node. If that is not the case, then naturally the shortest path between $s$ and $t$ is the edge $(s,t)$.
When $0<p_{i}\leq 1$ for all $1\leq i\leq m$, then $\log(1/p_{i})\geq 0$ for all $1\leq i\leq m$. Now, $\log(1/p_{i})$, for all $1\leq i\leq m$, can be used as a nonnegative distance between nodes, because as probability $p$ increases the distance $\log(1/p)$ decreases, i.e. $\log(1/p)$ is a strictly decreasing function of $p$. Furthermore, we can now use Dijkstra’s algorithm for finding shortest paths in the network, since edge weights are non-negative [10]. In this case, Dijkstra’s algorithm picks the edges with the lowest distance $\log(1/p_{i})$. Therefore, with this transformation the edge $i$ with the lowest value $\log(1/p_{i})$ is chosen i.e. the highest probability edge is chosen for the shortest path. In the special case, where $p_{i}=1$ for all $1\leq i\leq m$, then the problem reduces to finding the path with the fewest edges.